मराठी
महाराष्ट्र राज्य शिक्षण मंडळएचएससी विज्ञान (सामान्य) इयत्ता ११ वी

Solve the following: If P(A ∩ B) = 12, P(B ∩ C) = 13, P(C ∩ A) = 16 then find P(A), P(B) and P(C), If A,B,C are independent events.

Advertisements
Advertisements

प्रश्न

Solve the following:

If P(A ∩ B) = `1/2`, P(B ∩ C) = `1/3`, P(C ∩ A) = `1/6` then find P(A), P(B) and P(C), If A,B,C are independent events.

बेरीज
Advertisements

उत्तर

Since A and B are independent events,

P(A ∩ B) = P(A) · P(B)

∴ P(A) P(B) = `1/2`   ...(i)

B and C are independent events.

∴ P(B ∩ C) = P(B) · P(C)

∴ P(B) P(C) = `1/3`   ...(ii)

A and C are independent events.

∴ P(A ∩ C) = P(A) · P(C)

∴ P(A) P(C) = `1/6`  ...(iii)

Dividing (i) by (ii), we get

`("P"("A") * "P"("B"))/("P"("B") * "P"("C")) = (1/2)/(1/3)`

∴ P(A) = `3/2` P(C)   ...(iv)

Substituting equation (iv) in (iii), we get

`3/2`P(C) · P(C) = `1/6`

∴ [P(C)]2 = `1/9`

∴ P(C) = `1/3`

Substituting P(C) = `1/3` in equation (ii), we get P(B) = 1

Substituting P(B) = 1 in equation (i), we get P(A) = `1/2`

shaalaa.com
  या प्रश्नात किंवा उत्तरात काही त्रुटी आहे का?
पाठ 9: Probability - Miscellaneous Exercise 9 [पृष्ठ २१३]

APPEARS IN

बालभारती Mathematics and Statistics (Arts and Science) Part 1 [English] Standard 11 Maharashtra State Board
पाठ 9 Probability
Miscellaneous Exercise 9 | Q II. (4) | पृष्ठ २१३

व्हिडिओ ट्यूटोरियलVIEW ALL [2]

संबंधित प्रश्‍न

A die, whose faces are marked 1, 2, 3 in red and 4, 5, 6 in green is tossed. Let A be the event "number obtained is even" and B be the event "number obtained is red". Find if A and B are independent events.


Let E and F be events with `P(E) = 3/5, P(F) = 3/10 and P(E ∩ F) = 1/5`.  Are E and F independent?


Given that the events A and B are such that `P(A) = 1/2, PA∪B=3/5 and P (B) = p`. Find p if they are

  1. mutually exclusive
  2. independent.

Let A and B be independent events with P (A) = 0.3 and P (B) = 0.4. Find 

  1. P (A ∩ B)
  2. P (A ∪ B)
  3. P (A | B)
  4. P (B | A)

If A and B are two events such that `P(A) = 1/4, P(B) = 1/2 and P(A ∩ B) = 1/8`, find P (not A and not B).


Probability of solving specific problem independently by A and B are `1/2` and `1/3` respectively. If both try to solve the problem independently, find the probability that

  1. the problem is solved
  2. exactly one of them solves the problem.

One card is drawn at random from a well-shuffled deck of 52 cards. In which of the following case is the events E and F independent?

E : ‘the card drawn is a king or queen’

F : ‘the card drawn is a queen or jack’


If each element of a second order determinant is either zero or one, what is the probability that the value of the determinant is positive? (Assume that the individual entries of the determinant are chosen independently, each value being assumed with probability `1/2`).


If P(A) = 0·4, P(B) = p, P(A ⋃ B) = 0·6 and A and B are given to be independent events, find the value of 'p'.


A problem in statistics is given to three students A, B, and C. Their chances of solving the problem are `1/3`, `1/4`, and `1/5` respectively. If all of them try independently, what is the probability that, problem is solved?


A problem in statistics is given to three students A, B, and C. Their chances of solving the problem are `1/3`, `1/4`, and `1/5` respectively. If all of them try independently, what is the probability that, exactly two students solve the problem?


An urn contains four tickets marked with numbers 112, 121, 122, 222 and one ticket is drawn at random. Let Ai (i = 1, 2, 3) be the event that ith digit of the number of the ticket drawn is 1. Discuss the independence of the events A1, A2, and A3.


The odds against a husband who is 55 years old living till he is 75 is 8: 5 and it is 4: 3 against his wife who is now 48, living till she is 68. Find the probability that the couple will be alive 20 years hence.


The odds against student X solving a business statistics problem are 8: 6 and odds in favour of student Y solving the same problem are 14: 16 What is the chance that the problem will be solved, if they try independently?


A, B, and C try to hit a target simultaneously but independently. Their respective probabilities of hitting the target are `3/4, 1/2` and `5/8`. Find the probability that the target

  1. is hit exactly by one of them
  2. is not hit by any one of them
  3. is hit
  4. is exactly hit by two of them

The probability that a student X solves a problem in dynamics is `2/5` and the probability that student Y solves the same problem is `1/4`. What is the probability that

  1. the problem is not solved
  2. the problem is solved
  3. the problem is solved exactly by one of them

Two hundred patients who had either Eye surgery or Throat surgery were asked whether they were satisfied or unsatisfied regarding the result of their surgery

The follwoing table summarizes their response:

Surgery Satisfied Unsatisfied Total
Throat 70 25 95
Eye 90 15 105
Total 160 40 200

If one person from the 200 patients is selected at random, determine the probability that person was unsatisfied given that the person had eye surgery


Bag A contains 3 red and 2 white balls and bag B contains 2 red and 5 white balls. A bag is selected at random, a ball is drawn and put into the other bag, and then a ball is drawn from that bag. Find the probability that both the balls drawn are of same color


A family has two children. Find the probability that both the children are girls, given that atleast one of them is a girl.


If A and B′ are independent events then P(A′ ∪ B) = 1 – ______.


Three events A, B and C are said to be independent if P(A ∩ B ∩ C) = P(A) P(B) P(C).


For a loaded die, the probabilities of outcomes are given as under:
P(1) = P(2) = 0.2, P(3) = P(5) = P(6) = 0.1 and P(4) = 0.3. The die is thrown two times. Let A and B be the events, ‘same number each time’, and ‘a total score is 10 or more’, respectively. Determine whether or not A and B are independent.


A and B are two events such that P(A) = `1/2`, P(B) = `1/3` and P(A ∩ B) = `1/4`. Find: `"P"("B"/"A")`


A and B are two events such that P(A) = `1/2`, P(B) = `1/3` and P(A ∩ B) = `1/4`. Find: `"P"("A'"/"B'")`


If A and B are two events such that P(B) = `3/5`, P(A|B) = `1/2` and P(A ∪ B) = `4/5`, then P(A) equals ______.


If A and B are mutually exclusive events, then they will be independent also.


If A and B are two events such that P(A) > 0 and P(A) + P(B) >1, then P(B|A) ≥ `1 - ("P"("B'"))/("P"("A"))`


If A, B and C are three independent events such that P(A) = P(B) = P(C) = p, then P(At least two of A, B, C occur) = 3p2 – 2p3 


The probability of obtaining an even prime number on each die when a pair of dice is rolled is


Two events 'A' and 'B' are said to be independent if


Let A and B be independent events P(A) = 0.3 and P(B) = 0.4. Find P(A ∩ B)


Let EC denote the complement of an event E. Let E1, E2 and E3 be any pairwise independent events with P(E1) > 0 and P(E1 ∩ E2 ∩ E3) = 0. Then `"P"(("E"_2^"C"  ∩ "E"_3^"C")/"E"_1)` is equal to ______.


Five fair coins are tossed simultaneously. The probability of the events that at least one head comes up is ______.


A die is thrown once. What is the sample space \[S\]?


A die is thrown once. If F is the event that the number appearing is even, which set represents F?


For \[E=\{3,6\}\] when a die is thrown once, what is \[P(E)\]?


For \[P(E)=\frac{1}{3}\] and \[P(F)=\frac{1}{2}\], what is \[P(E)P(F)\]?


Share
Notifications

Englishहिंदीमराठी


      Forgot password?
Use app×